a) Bara enterprises operates in Kitale town. On annual basis, it orders 480,000 pens form a...

  1. a) Bara enterprises operates in Kitale town. On annual basis, it orders 480,000 pens form a Nairobi based distributor. A packet of 24 pens delivered to Baras warehouse cost sh 480 including transport charges. The supermarket barrows money from ABC bank of an interest rate of 10% p.a to finance its inventories. The supermarket also incurs sh 1,500 to place an order for the pens and sh 8 carrying cost for each pen

REQUIRED:

  1. Calculate the economic order quantity (EOQ) for the pens
  2. Total cost of the EOQ
  3. For orders for 72,000 pens and above the distributer has offered a discount rate of 10% on delivery price.

Advice the management of the supermarket on whether to take advantages of the discount offer.

b) Wilson an automotive technician has been operating a garage in Mombasa for the past two years. A year ago he converted part of his garage to a welding shop making and selling metal doors and windows. He had anticipated that he cost of welding shop would primarily be final but has realized that the welding cost increased with the increase in number of welding job assignment. The costs of welding job assignments are as follows.

Perched

No of welding job assignments

Total cost sh ‘000’

Sept 2008

   280

700

Oct   2008

   800

860

Nov 2008

1,240

110

Dec   2008

1,000

960

Jan   2009

   600

720

Feb   2009

   920

910

Mar 2009

   860

880

Apr   2009

1,200

260

REQUIRED:

Formulate an equation to estimate the total cost of the welding shop and compute the cost of undertakings 1256 assignment using

  1. High-low method
  2. Simple linear regression method.                               (6 Marks)
Answer #1

Dear student, we cannot able to post solution more than one question as per our policy.

Part 1

Part 2

Purchasing cost (480000*20)          9,600,000
Ordering cost (Number of order = 480000/12000=40) (40*1500)                60,000
Carrying cost (Average stock = 12000/2 = 6000) (6000*10)                60,000
Total annual cost under the EOQ Policy          9,720,000

Part 3

New purchase price per pen (20-(20*10%))                        18
Carrying cost per pen                     8.00
Add: annual finance cost per pen (18*10%)                     1.80
Annul carrying cost per pen (C)                     9.80
Purchasing cost (480000*18)          8,640,000
Ordering cost (Number of order = 480000/72000=6.67 = 7 Orders) (7*1500)                10,500
Carrying cost (Average stock = 72000/2 = 36000) (36000*9.80)              352,800
Total annual cost under using the discount offer          9,003,300
Advice the management of the supermarket on whether to take advantages of the discount offer.
Total annual cost under the EOQ Policy          9,720,000
Total annual cost under using the discount offer          9,003,300
Lowest cost is better option.
Management should take advantages of the discount offer.